OpenStudy (kaynick):

The points (−1,r) and (4,4) lie on a line with slope −2. Find the missing coordinate r.

OpenStudy (greatlife44):

interesting

OpenStudy (greatlife44):

We can use this equation here \[y-y_{1} = m(x-x_{1})\] r = y1 then we solve for y1 \[y_{1} = y-m(x-x_{1})\] plug in everything \[4-(-2)*(4-(-1)) = 4-(-2*5) = 14\] this means that y_{1} = 14 = r now let's check this (-1,14) (4,4) \[m = \frac{ y-y_{1} }{ x-x_{1} } = \frac{ 14-4 }{ -1-4 } = \frac{ 10 }{ -5 } = -2 \]

OpenStudy (anonymous):

Refer to the Mathematica attachment.